OpenStudy (anonymous):

find the constant variation k. y varies directly with x; y = 750 when x =1500 k = ?

Directrix (directrix):

1/2 -- need to check

OpenStudy (anonymous):

y = kx -> k = y/x = 750/ 1500 = 1/2